# Golden Touchstone: A Comprehensive Bilingual Benchmark for Evaluating Financial Large Language Models

### v1

@@ Abstract
As large language models become increasingly prevalent in the financial sector, there is a pressing need for a standardized method to comprehensively assess their performance. However, existing finance benchmarks often suffer from limited language and task coverage, as well as challenges such as low-quality datasets and inadequate adaptability for LLM evaluation. To address these limitations, we propose "Golden Touchstone", the first comprehensive bilingual benchmark for financial LLMs, which incorporates representative datasets from both Chinese and English across eight core financial NLP tasks. Developed from extensive open source data collection and industry-specific demands, this benchmark includes a variety of financial tasks aimed at thoroughly assessing models’ language understanding and generation capabilities. Through comparative analysis of major models on the benchmark, such as GPT-4o, Llama3, FinGPT and FinMA, we reveal their strengths and limitations in processing complex financial information. Additionally, we open-sourced Touchstone-GPT, a financial LLM trained through continual pre-training and financial instruction tuning, which demonstrates strong performance on the bilingual benchmark but still has limitations in specific tasks.This research not only provides the financial large language models with a practical evaluation tool but also guides the development and optimization of future research. The source code for Golden Touchstone and model weight of Touchstone-GPT have been made publicly available at https://github.com/IDEA-FinAI/Golden-Touchstone, contributing to the ongoing evolution of FinLLMs and fostering further research in this critical area.
{CJK*}
UTF8gbsn
11footnotetext: Equal Contribution.22footnotetext: Work done during internship at IDEA Research.33footnotetext: Corresponding Author.footnotetext: Correspondence to: Jian Guo <guojian@idea.edu.cn>.
@@ 1 Introduction
The rapid development of both proprietary(Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Anthropic, 2024; Team et al., 2023) and open-source Large Language Models (LLMs) has led to their increasing importance and application across various fields(Touvron et al., 2023a, b; AI@Meta, 2024; Bai et al., 2023; Yang et al., 2024a; DeepSeek-AI, 2024; Young et al., 2024; Zeng et al., 2023; Baichuan, 2023; Gan et al., 2023; Zhang et al., 2022) , including finance(Wu et al., 2023; Lopez-Lira and Tang, 2023), healthcare Thirunavukarasu et al. (2023); Tian et al. (2023), and law Cui et al. (2023); Xiao et al. (2021). Among these, the financial sector stands out as one of the most important areas for LLM application, due to its rich textual information and high practical value.
In recent years, a variety of advanced financial large language models (FinLLMs) have emerged, capable of specialized tasks such as financial sentiment analysis, content summarization, stock movement prediction, and specialized question answering, as depicted in Figure.1. These models, including FinGPT (Yang et al., 2023), PIXIU Xie et al. (2023), CFGPT (Li et al., 2023), DISC-FinLLM (Chen et al., 2023), and XuanYuan (Zhang and Yang, 2023), leverage unique frameworks and tuning methods to enhance their performance on domain-specific benchmarks, offering robust solutions for real-world financial applications. Alongside the rapid emergence of FinLLMs, there has also been a significant increase in financial benchmarks. In the English financial domain, FLUE(Shah et al., 2022) was the first publicly available benchmark for assessing English financial language understanding. Similarly, BBT(Lu et al., 2023a) was the pioneering open-source financial NLP benchmark in Chinese. As LLMs and financial applications have continued to develop, more financial benchmarks have been introduced(Xie et al., 2023, 2024; Yang et al., 2023; Lei et al., 2023; Zhang et al., 2023). While these benchmarks have provided valuable resources and evaluation criteria for assessing current LLMs in finance, there remain challenges, such as the presence of low-quality or unsuitable for LLMs datasets within certain tasks, leading to generally poor evaluation results across various models. Furthermore, Current benchmarks also suffer from insufficient language and task coverage, which prevents a comprehensive evaluation of large financial language models.
@@
To address these gaps, we propose Golden Touchstone, a bilingual financial benchmark that consolidates representative datasets across eight financial NLP tasks in both Chinese and English. Golden Touchstone provides high-quality datasets, task-aligned metrics, and instructional templates tailored to guide LLMs in generating task-appropriate responses. The Golden Touchstone organizes all task data into an instruction, input, and output format, facilitating the use and evaluation of different models. We evaluated current open-source general-purpose LLMs and FinLLMs using the Golden Touchstone benchmark. Results indicate that while GPT-4o, Qwen-2, Llama-3, and FinMA performed well across several tasks like financial sentiment analysis and entity extration, there remains considerable room for improvement in areas like credit card scoring dataset within classification tasks and stock movement prediction. Additionally, we have open-sourced Touchstone-GPT, a model trained through domain-specific continual pre-training and financial instruction tuning. It achieved acceptable performance on the bilingual benchmark but showed limitations in tasks such as stock movement prediction and question answering. These findings underscore the need for more high-quality training data and potentially more suitable model architectures for finance-specific applications. Our main contributions are as follows:
@@
  • Introduction of Golden Touchstone, the first comprehensive bilingual benchmark for financial Large Language Models (LLMs), encompassing 22 datasets across eight tasks in both Chinese and English. Compared to previous FinLLMs’ benchmarks, our benchmark offers enhanced diversity, systematicity, and adaptability.
  • A thorough evaluation of state-of-the-art LLMs and FinLLMs, including GPT-4o, Qwen-2, Llama-3, FinGPT and FinMA and so on, on the Golden Touchstone. This evaluation highlights the key strengths and shortcomings in model performance across various tasks, and suggests directions for future research in financial LLMs.
  • The open-source release of Touchstone-GPT, a specialized financial LLM that has undergone domain-specific continuous pre-training and instruction tuning. This model serves as a new baseline for subsequent FinLLMs research, fostering further advancements in Financial AI.

### v2

@@ Abstract
As large language models (LLMs) increasingly permeate the financial sector, there is a pressing need for a standardized method to comprehensively assess their performance. Existing financial benchmarks often suffer from limited language and task coverage, low-quality datasets, and inadequate adaptability for LLM evaluation. To address these limitations, we introduce Golden Touchstone, a comprehensive bilingual benchmark for financial LLMs, encompassing eight core financial NLP tasks in both Chinese and English. Developed from extensive open-source data collection and industry-specific demands, this benchmark thoroughly assesses models’ language understanding and generation capabilities. Through comparative analysis of major models such as GPT-4o, Llama3, FinGPT, and FinMA, we reveal their strengths and limitations in processing complex financial information. Additionally, we open-source Touchstone-GPT, a financial LLM trained through continual pre-training and instruction tuning, which demonstrates strong performance on the bilingual benchmark but still has limitations in specific tasks. This research provides a practical evaluation tool for financial LLMs and guides future development and optimization. The source code for Golden Touchstone and model weight of Touchstone-GPT have been made publicly available at https://github.com/IDEA-FinAI/Golden-Touchstone.
11footnotetext: Equal Contribution: wuxiaojun@idea.edu.cn, liujunxi@m.scnu.edu.cn22footnotetext: Corresponding Author: jialee@connect.hkust-gz.edu.cn, guojian@idea.edu.cn
@@ 1 Introduction
The rapid development of both proprietary (brown2020gpt3; ouyang2022instruct_gpt; openai2023gpt4; anthropic2024claude; team2023gemini) and open-source Large Language Models (LLMs) (touvron2023llama; touvron2023llama2; llama3; bai2023qwen; yang2024qwen2; deepseekv2; young2024yi; zeng2023glm-130b; baichuan2023baichuan2; gan2023ziya2; zhang2022fengshenbang) has led to their increasing application in various fields, including finance (wu2023bloomberggpt; lopez2023chatgpt_stock_move), healthcare (thirunavukarasu2023llm_med; tian2023chimed), and law (cui2023chatlaw; xiao2021lawformer). Among these, the financial sector shown in Figure.1 stands out as a critical area for LLM application due to its rich textual information and high practical value.
In recent years, a variety of advanced financial large language models (FinLLMs) have emerged, capable of specialized tasks such as financial sentiment analysis, content summarization, stock movement prediction, and question answering (yang2023fingpt; xie2023pixiu; li2023cfgpt; chen2023disc; zhang2023xuanyuan). These models leverage unique frameworks and tuning methods to enhance their performance on domain-specific benchmarks, offering robust solutions for real-world financial applications. However, existing financial benchmarks often suffer from limited language and task coverage, low-quality datasets, and inadequate adaptability for LLM evaluation, leading to poor evaluation results (shah2022flue; lu2023bbt; xie2023pixiu; xie2024finben; yang2023fingpt; lei2023cfbenchmark; zhang2023fineval).
@@
To address these challenges, we propose Golden Touchstone, a comprehensive bilingual benchmark for financial LLMs, encompassing eight core financial NLP tasks in both Chinese and English. Golden Touchstone provides high-quality datasets, task-aligned metrics, and instructional templates to guide LLMs in generating task-appropriate responses. We evaluated several state-of-the-art models, including GPT-4o, Qwen-3, Llama-3, FinGPT, and FinMA, on this benchmark. Results indicate that while these models perform well on tasks such as sentiment analysis and entity extraction, there is significant room for improvement in areas like stock movement prediction and classification tasks. Additionally, we open-source Touchstone-GPT, a financial LLM trained through domain-specific continual pre-training and instruction tuning, which serves as a new baseline for future research.
Our contributions are as follows:
@@
  • Introduction of Golden Touchstone, a comprehensive bilingual benchmark for financial LLMs, encompassing 22 datasets across eight tasks in both Chinese and English.
  • Evaluation of state-of-the-art LLMs and FinLLMs on Golden Touchstone, highlighting their strengths and limitations across various tasks.
  • Open-sourcing of Touchstone-GPT, a financial LLM trained through domain-specific continual pre-training and instruction tuning, fostering further advancements in financial AI.
2 Benchmark Design
2.1 Current Benchmark Status
Existing open-source financial benchmarks have made significant strides in evaluating financial natural language processing (NLP) tasks. FLUE (shah2022flue) pioneered English financial NLP evaluation, covering sentiment analysis, news classification, and other critical tasks. Subsequently, PIXIU (xie2023pixiu) and FinBen (xie2024finben) expanded task coverage, while in the Chinese domain, BBT-Benchmark (lu2023bbt) introduced the first comprehensive Chinese financial evaluation framework.
These benchmarks suffer from several critical limitations. Firstly, there is inconsistent data quality across different tasks, as shown in Table 1. Additionally, large language models face challenges in numerical understanding, a problem highlighted in various studies (shen2023positional; akhtar2023numerical_reasoning; schwartz2024numerologic; hu2024no). Lastly, the benchmarks lack bilingual assessment capabilities, a gap that has been pointed out in recent research (nie2024cfinbench; lei2023cfbenchmark). These limitations collectively hinder the overall effectiveness and reliability of the benchmarks. To address these limitations, we propose a unified benchmark that integrates high-quality financial datasets from both English and Chinese domains. Our approach aims to provide a more comprehensive and linguistically diverse evaluation of financial large language models (LLMs).
Table 1: Diversity of Financial Analysis Tasks Across Different Benchmarks
FinGPT-Bench (wang2023fingptbench) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FinBen (xie2024finben) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BBT-Fin (lu2023bbt) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Fin-Eval (zhang2023fineval) $\checkmark$
CFBenchmark (lei2023cfbenchmark) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2.2 Golden Touchstone Benchmark Framework
Addressing these critical limitations, we introduce the Golden Touchstone benchmark, a comprehensive bilingual evaluation framework designed to holistically assess financial language models. Conceptualized around two primary dimensions—task types and language coverage—our approach represents a significant departure from existing evaluation methodologies. The overview framework can be seen in Figure.2. Firstly, our Golden Touchstone benchmark categorizes financial NLP tasks across two dimensions: the first dimension is task type including Natural Language Understanding (NLU) including sentiment analysis, knowledge examination, information extraction, text understanding tasks and Natural Language Generation (NLG) including summary analysis, question answering tasks, and the second dimension is language including English and Chinese.
The benchmark strategically integrates English and Chinese datasets across eight sophisticated sub-tasks, spanning Natural Language Understanding (NLU) and Natural Language Generation (NLG). By carefully curating high-quality datasets from existing benchmarks and datasets. The detailed data statistics are shown in Appendix.A. The benchmark encompasses eight critical sub-tasks:
  • Sentiment Analysis: Utilizing datasets like FPB (malo2014fpb) and FiQA-SA (maia2018finqa_sa) and FinFE-CN (lu2023bbt)
  • Classification: Integrating Headlines (sinha2021headlines), FOMC (shah2023fomc), and LendingClub (feng2023lendingclub) and FinNL-CN (lu2023bbt) datasets
  • Entity Recognition: Using NER (alvarado2015NER) and FinESE-CN (lu2023bbt) datasets
  • Relation Extraction: Incorporating FinRED (sharma2022finred) and FinRE-CN (lu2023bbt) datasets
  • Multiple Choice: Drawing from CFA (yang2024ideafinbench) and FinEval (zhang2023fineval) and CPA (yang2024ideafinbench) datasets
  • Summarization: Employing EDTSUM (zhou2021edtsum) and FinNA-CN lu2023bbt datasets
  • Question Answering: Utilizing FinQA (chen2021finqa) and FinQA-CN and FinCQA-CN (lu2023bbt) datasets
  • Stock Movement Prediction: Introducing news-based prediction using CMIN-US and CMIN-CN (luo2023cmin) datasets
A implementation consideration involves the treatment of [unknown] labels across various tasks. Our approach consistently treats [unknown] labels as incorrect predictions, which maintains uniform evaluation standards across all tasks, avoids potential biases introduced by alternative methods, and reflects real-world scenarios where models must make definitive predictions. This approach proves superior to alternative methods such as sample omission, random assignment, or fixed selection, which may introduce evaluation biases or distort performance metrics. The benchmark also design incorporates several features to ensure meaningful cross-lingual comparisons, including maintaining identical task structures across English and Chinese benchmarks, using comparable evaluation metrics for equivalent tasks, and choosing datasets with similar characteristics and difficulty levels. These design choices enable robust comparison of model performance across languages while accounting for linguistic and cultural differences in financial contexts.
By addressing previous benchmarks’ limitations, our approach provides a more robust, comprehensive, and linguistically diverse framework for evaluating financial large language models. The benchmark not only expands task coverage but also addresses critical challenges in current financial NLP evaluation methodologies. Our key methodological innovations reflect significant efforts in data selection, cleaning, and organization. We replaced time-series tabular data with news-based stock prediction to enhance the relevance and diversity of the dataset. Additionally, we ensured bilingual task alignment to address the lack of multilingual assessment capabilities, while carefully selecting datasets with consistent and high-quality labeling to overcome issues related to data quality. Furthermore, we developed a unified evaluation approach across different financial NLP tasks to ensure robust and comparable results. In line with our commitment to fostering research transparency and collaboration, we have not only open-sourced the benchmark datasets but also released a well-documented, clear, reproducible, and extensible evaluation code framework for the broader research community to utilize and build upon. The Golden Touchstone benchmark represents a significant step forward in financial NLP evaluation, addressing key limitations of existing benchmarks while providing a comprehensive, bilingual framework for assessing model performance. Its systematic design, rigorous evaluation methodology, and practical focus make it an essential tool for researchers and practitioners in financial NLP.
3 Experiments
3.1 Experimental Setup
Baselines. We conducted an extensive experimental evaluation against the Golden Touchstone Benchmark, incorporating a comprehensive array of models. For all models and inference tasks, we set the PyTorch and CUDA random seeds and configured the model with a greedy decoding strategy. This ensures reproducibility of experimental results and eliminates the influence of sampling decoding strategies on the final generated outputs. This included cutting-edge commercial models such as GPT-4o (openai2023gpt4), alongside prominent open-source alternatives like Meta Llama-3 (llama3) and Alibaba Qwen-3 (yang2025qwen3). Additionally, we integrated the latest and most influential financial language models (FInLLMs), namely FinGPT (yang2023fingpt), FinMA (xie2024finben), CFGPT (li2023cfgpt), and DISC-FinLLM chen2023disc. These models were meticulously selected to represent a diverse spectrum of capabilities, ranging from general-purpose language understanding to specialized financial domain expertise. Our experiments aimed to rigorously assess the performance, robustness, and adaptability of each model within the context of financial data processing and analysis. The results provide valuable insights into the strengths and limitations of current state-of-the-art models, offering a foundation for future advancements in financial language modeling.
Touchstone-GPT Training. To further contribute to the research and development of FInLLMs and Financial benchmarks for LLMs, we have meticulously trained and open-sourced a Touchstone-GPT model. This initiative aims to serve as a valuable resource for advancing the field, providing a robust and versatile model that can be utilized for a wide range of financial language tasks. We adopted a two-stage training strategy comprising continuous pre-training and post-training, based on the Qwen-2.5 (yang2024qwen2) foundational model. During the continuous pre-training phase, we initially conducted pre-training on a high-quality financial corpus containing 100 billions tokens, which included textbooks, encyclopedias, research reports, news articles, and real-time analysis, all meticulously cleaned. In the post-training phase, we employed a standard instruction fine-tuning strategy, collecting, cleaning, and formatting a high-quality dataset of 300,000 instruction-response pairs shown in Tabel.4 and Table.5. To avoid catastrophic forgetting in general tasks, we also incorporated general-domain pre-training corpora (gan2023ziya2) and instruction-tuning corpora (peng2023Alpaca_gpt4) into continuous pre-training and post-training. This culminated in the final Touchstone-GPT model. We utilized Megatron(shoeybi2019megatron) for continuous pre-training and LlamaFactory(zheng2024llamafactory) for instruction post-training as our training frameworks, respectively. In this study, we employ an advanced model training setup using the AdamW optimizer (kingma2014adam) with a learning rate of 1.0e-5, cosine annealing scheduler (szegedy2016cosine), and a 10% warmup ratio (he2016resnet; goyal2017warmup) to enhance training stability and convergence. We enable gradient accumulation (shoeybi2019megatron) and checkpointing (chen2016gradient_ckpt) to simulate larger batch sizes and reduce memory footprint. Training is conducted in mixed bfloat16 precision (micikevicius2017mixed; wang2019bf16) with DeepSpeed’s ZeRO-1 optimization (rajbhandari2020zero), reducing memory consumption and allowing for larger model training. This comprehensive setup optimizes efficiency and performance, providing an effective solution for large-scale deep learning model training.
Our training was conducted on 4 NVIDIA DGX servers, each equipped with 8 A100 GPUs, and spanned a period of 4 weeks. Inference was performed on a single NVIDIA DGX server with eight A100 GPUs, utilizing parallel batch inference. During the pre-training phase, we employed a data packing strategy (krell2021packing) and batch dynamic right padding strategy in the instruction tuning phase (wolf-etal-2020-transformers), while the inference phase incorporated a batch left padding strategy (wolf-etal-2020-transformers).
3.2 Evaluation Results
From the perspective of individual models in Figure.3, GPT-4o shows strong performance in sentiment analysis and structured tasks like multiple choice, indicating robust general language understanding capabilities. However, its weakness lies in relation extraction and detailed entity extraction, which require detailed understanding of complex financial relations. FinMA-7B stands out in sentiment tasks but lacks versatility, especially in question answering and summarization, likely due to the absence of targeted training for diverse NLP challenges. Qwen-3-8B demonstrates a balanced yet still improvable performance profile: it achieves decent results in sentiment analysis tasks but shows notable deficiencies in more complex tasks like question answering and summarization, indicating the need for more targeted domain adaptation training to enhance its specialized capabilities. Llama-3-8B-Instruct excels in english NLU tasks, but shows limitations in tasks requiring chinese tasks, such as entity and relation analysis. The metrics of FinGPT-8B-lora indicating that the current level of domain-specific tuning is insufficient for complex financial tasks. Finally, DISC-FinLLM-Full and CFGPT1-7B-Full demonstrate moderate strengths in entity extraction tasks but lack the robustness needed for broader NLP capabilities, revealing significant gaps in financial language comprehension.
Figure 3: Comparison of performance from the perspective of models. Each subplot represents the performance of a models on both English and Chinese tasks. The bars indicate the model’s performance on each task, while the dashed gray line represents the average performance across all models for that task.
From a task perspective in Figure.4, we observe that Sentiment Analysis generally yields high scores across most models, particularly for the English benchmark, indicating that sentiment understanding, even in financial contexts, is relatively well addressed by these models. In contrast, Relation Extraction and Question Answer in financial domain exhibit notably lower performance, especially for the Chinese benchmark. These results suggest that capturing financial relationships and classifying detailed financial statements pose greater challenges, requiring more sophisticated training datasets or better model architectures. The LendingClub dataset in Classification is a specialized dataset in the field of risk control, requiring more targeted fine-tuning to achieve good results. Stock Movement Prediction also shows low performance across most models, with only a few models such as GPT-4o demonstrating relatively moderate performance, but it is still practically unusable, highlighting the inherent difficulty of this task. Market prediction relying solely on news information is likely insufficient; volume-price data and factor analysis can provide more comprehensive information. However, current large language models are unable to process these inputs, which is a significant area of future research. Summarization also stands out as a weak area for most models, with consistently low BLEU and Rouge scores, reflecting the challenges in generating concise, coherent summaries of complex financial text.
Figure 4: Comparison of performance from the perspective of tasks, illustrating average performance for English and Chinese tasks respectively.
Table 2: Performance metrics of financial large language models across English tasks like Sentiment Analysis, Classification, and Summarization. Models include GPT-4o, Llama-3-8B, Qwen-3-8B, FinMA-7B, FinGPT-8B, Mistral-7B, and Touchstone-GPT. The best results of each dataset are marked in bold.
Task Dataset Metrics GPT-4o FinMA-7B full Qwen-3-8B Llama-3-8B Instruct FinGPT-8B lora Mistral-7B Touchstone GPT
Sentiment
Analysis
FPB Weighted-F1 0.8084 0.9400 0.8100 0.7631 0.2727 0.8110 0.8576
ACC 0.8093 0.9402 0.8020 0.7660 0.3072 0.8030 0.8557
Fiqa-SA Weighted-F1 0.8106 0.8370 0.6800 0.7515 0.5885 0.6810 0.8591
ACC 0.7702 0.8340 0.6000 0.7064 0.5872 0.6010 0.8638
Classification
Headlines Weighted-F1 0.7857 0.9739 0.7300 0.7006 0.4516 0.7310 0.9866
ACC 0.7931 0.9739 0.7270 0.7004 0.4331 0.7280 0.9866
FOMC Weighted-F1 0.6603 0.3988 0.6150 0.4904 0.2758 0.6160 0.8788
ACC 0.6794 0.4274 0.6230 0.5625 0.2702 0.6240 0.8790
lendingclub Weighted-F1 0.6730 0.1477 0.5960 0.5943 0.5480 0.5970 0.9783
MCC 0.1642 -0.6218 0.1720 0.1670 -0.1120 0.1730 0.9297
Entity
Extraction
NER Entity-F1 0.1800 0.6200 0.2900 0.2973 0.0231 0.2910 0.6993
Relation
Extraction
FinRE Relation-F1 0.1613 0.0054 0.1100 0.0540 0.0100 0.1110 0.5331
Multiple
Choice
CFA Weighted-F1 0.7700 0.2200 0.6720 0.5800 0.3993 0.6730 0.7497
ACC 0.7700 0.2400 0.6720 0.5800 0.3800 0.6730 0.7500
Summarization
EDTSUM Rouge-1 0.1675 0.1566 0.1480 0.1467 0.0622 0.1490 0.5254
Rouge-2 0.0556 0.0491 0.0440 0.0429 0.0085 0.0450 0.3446
Rouge-L 0.1069 0.1060 0.0860 0.0930 0.0412 0.0870 0.4705
BLEU 0.1192 0.1361 0.1010 0.1085 0.0592 0.1020 0.4512
Question
Answering
Finqa RMACC 0.1037 0.0497 0.0280 0.0470 0.0110 0.0290 0.2258
Convfinqa RMACC 0.2540 0.0953 0.0650 0.1477 0.0772 0.0660 0.5053
Stock Movement
Prediction
CMIN-US Weighted-F1 0.5025 0.2639 0.4130 0.3722 0.3379 0.4140 0.5036
ACC 0.5149 0.3446 0.5120 0.4955 0.4154 0.5130 0.5144
Table 3: Performance metrics of financial large language models across chinese tasks like Sentiment Analysis, Classification, and Summarization. Models include GPT-4o, Llama-3-8B, Qwen-3-8B, CFGPT-7B, DISC-FinLLM, and Touchstone-GPT. The best results of each dataset are marked in bold.
Task Dataset Metrics GPT-4o Qwen-3-8B Instruct Llama-3-8B Instruct CFGPT1-7B Full DISC-FinLLM Full Touchstone GPT
Sentiment
Analysis
FinFe-CN Weighted-F1 0.6593 0.6478 0.3633 0.2528 0.4177 0.7888
ACC 0.6500 0.6632 0.4891 0.2732 0.4292 0.7936
Classification
FinNL-CN ORMACC 0.3303 0.0824 0.0747 0.0894 0.0011 0.8360
Entity
Extraction
FinESE-CN ORMACC 0.6867 0.3879 0.3088 0.3863 0.4346 0.9074
Relation
Extraction
FinRE-CN RMACC 0.2754 0.1532 0.1296 0.0678 0.1182 0.6541
Multiple
Choice
FinEval Weighted-F1 0.7364 0.7332 0.4432 0.3543 0.4288 0.7361
ACC 0.7353 0.7337 0.4471 0.3529 0.4294 0.7353
CPA Weighted-F1 0.6312 0.7159 0.3421 0.3543 0.3451 0.9238
ACC 0.6309 0.7162 0.3504 0.3553 0.3518 0.9238
Summarization
FinNA-CN Rouge-1 0.3197 0.3528 0.3477 0.1018 0.3486 0.5526
Rouge-2 0.1434 0.1699 0.1702 0.0263 0.1678 0.3603
Rouge-L 0.2511 0.2846 0.2802 0.0650 0.2997 0.5214
BLEU 0.1423 0.1743 0.1672 0.0238 0.1885 0.3944
Question
Answering
FinQa-CN RMACC 0.6578 0.5245 0.4540 0.1126 0.3949 0.9214
FinCQa-CN RMACC 0.4765 0.3624 0.3787 0.2714 0.2134 0.8552
Stock Movement
Prediction
CMIN-CN Weighted-F1 0.4858 0.4165 0.4497 0.3549 0.0329 0.4735
ACC 0.4988 0.4825 0.4858 0.3584 0.0332 0.4878